OpenStudy (anonymous):

if 2x=12 which of the folwing is true 2x-7=12-7 2x-7=12+7

OpenStudy (anonymous):

When we solve algebraic equations like this it can help to think of the equation like a teeter-totter or old scale balanced on the equals sign. In order to keep it balanced we have to make sure that whatever we do to one side of the equals sign, we do the exact same thing to the other side. In the first case we subtracted 7 from one side so we need to make sure we subtract 7 from the other side as well. Since this is the case, this equation is true. In the second case we subtracted 7 from one side so we need to make sure we subtract 7 from the other side as well. In this case the right hand side has a +7, not a -7 to match the left hand side, so this equation is false. Answer: 2x-7=12-7 is TRUE. 2x-7=12+7 is FALSE.